OpenStudy (anonymous):

A cube has v vertices, e edges, and f faces. What is the value of f × v – e?

OpenStudy (anonymous):

A cube normally has 8 vertices, 12 edges, and 6 faces so therefore: \[f\times v-e\] \[\eqalign{ &f=6 \\ &v=8 \\ &e=12 \\ }\] So therefore: \[f\times v-e=6\times8-12=48-12=36\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."